# PlanScope: Learning to Plan Within Decision Scope Does Matter

### v1

@@ III-B Iterative Detail Decoder
$\frac{1}{N+2}$ $\left(\sum^{N}_{l=0}\left(\lVert\mathcal{T}_{D}^{l}[:H_{l}]-\hat{\mathcal{T}}_{D}^{l}[:H_{l}]\rVert_{2}\right)+\lVert\mathcal{T}_{A}^{N}-\hat{\mathcal{T}}_{A}^{N}\rVert_{2}\right),$
where $N$ is the count of detail level. The final query $Q_{N+1}$ is then forwarded to the trajectory MLP decoder and scoring MLP decoder, yielding the multi-modal trajectory and the score by
$\mathcal{T}=\operatorname{MLP}(Q_{N+1})$ $,\,\pi=\operatorname{MLP}(Q_{N+1}),$
$\mathcal{L}_{\text{reg}}=\operatorname{L1_{smooth}}\left(\mathcal{T},\hat{\mathcal{T}}\right)$ $,\,\mathcal{L}_{\text{cls}}=\operatorname{CrossEntropy}\left(\pi,\hat{\pi}\right),$
$\mathcal{L}=\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{cls}}+\mathcal{L}_{\text{pre}}+\mathcal{L}_{\text{col}}+\mathcal{L}_{\text{ds}}.$
@@ IV Experiments
Figure 5: In scenario S1, PLUTO plans to merge into the right lane, but it extensively focuses on future plans and does not pay attention to the vehicle approaching from the right rear, which causes the collision when T=10. PlanScope focused on the right rear and briefly made an evasive maneuver. In scenario S2, a typical left turn at an unprotected intersection, PLUTO fails to capture the sudden acceleration operation in the log replay, fearing the agent driving from opposite stopping in lane. This decision result in missing the action window and being rear-ended by the following vehicle, while PlanScope avoids the accident.
@@ IV-A Implementation Details
Our training is carried out in the nuPlan dataset [5], which comprises a training set of 25,000 scenarios. We evaluate our model’s performance against baseline models within the nuPlan challenge framework. In addition, we conduct an ablation study to determine the significance of each component within our model. The training is performed on a server equipped with four NVIDIA RTX 3090 GPUs. We employ the Adam optimizer, with a start learning rate of 0.0001, 3 warm-up epochs, and a batch size of 32, over the course of 25 training epochs. The PyTorch library is utilized for the model’s implementation, with each training session approximately lasting 2 days.
@@
Our evaluation is conducted on the Val14 [25] subset of the nuPlan dataset, encompassing 1,090 validating scenarios. Each simulation entails a 15-second rollout at a frequency of 10 Hz. The evaluation score is derived using the official evaluation script provided by the nuPlan challenge. Our primary comparison metric is the Non-Reactive Closed-Loop Simulation (CLS-NR) driving score, which is more challenging as other vehicles do not avoid collisions with the ego vehicle. Besides, the future steps of our model is 80 while the history steps is 21. Both our model and PLUTO [21] are set to generate 6 modal trajectories. In our experiments, the maximum horizons of decision scope across levels are controlled by setting DS Horizon, which is denoted as $h$. Furthermore, we borrow the idea in [26] to facilitate the rapid implementation and execution of wavelet transformation in our experiments.
IV-B Comparison with Baselines
TABLE I: Comparison with Baseline Methods
Planner Comfortness Progress Collisions Speed Limit Drivable TTC Score
Log-Replay 99.27% 98.99% 98.76% 96.47% 98.07% 94.40% 93.68%
IDM [27] 79.31% 86.16% 90.92% 97.33% 94.04% 83.49% 79.31%
RasterModel [5] 81.64% 80.60% 86.97% 98.03% 85.04% 81.46% 66.92%
UrbanDriver [28] 100.00% 80.83% 94.13% 91.58% 90.83% 80.28% 67.72%
PLUTO [21] 95.14% 87.13% 94.45% 97.87% 98.90% 90.92% 85.81%
PlanScope-h10 (ours) 96.24% 87.59% 94.59% 98.03% 98.99% 91.10% 85.94%
PlanScope-h20 (ours) 94.13% 85.39% 95.41% 98.48% 99.72% 93.94% 85.97%
@@ IV-B Comparison with Baselines
We conduct a comparative analysis of our method against several standard approaches, the comparative results are presented in Table I. Log-Replay, which entails a straightforward replay of expert trajectories, serves as a baseline reference. This demonstrates the capability of our simulator to accurately assess the performance of planning algorithms. IDM [27] employs the Intelligent Driver Model to maintain safe distances from other agents and adhere to the reference lane. RasterModel, introduced in [5], is a Convolutional Neural Network (CNN)-based planning model. UrbanDriver [28] is a vectorized planning model leveraging PointNet based polyline encoders and transformer architecture. PLUTO [21], our primary baseline, achieves state-of-the-art performance in the nuPlan planning challenge in April 2024. Naturally, we have excluded the rule-based trajectory selector, thereby demonstrating that our new model enhances the performance of the neural planner. Besides, we also eliminate the contrastive learning module, which is empowered by rule-based positive and negative sample construction. Our model, designated as PlanScope, is evaluated with maximum horizons of 10 and 20, decomposing velocity profile, and the maximum decomposition level $N$ across experiments is 3. A qualitative comparison is depicted and introduced in Fig. 5. The results, as depicted in the table, indicate that the optimal total score of 85.97% is achieved at a horizon of 20. Notably, key safety metrics, including Time To Collision (TTC) and the absence of Ego-at-fault Collisions (Collisions), exhibit significant enhancements of 3.02% and 0.96%, respectively. At a horizon of 10, although the total score is marginally reduced, our model still surpasses the baseline across nearly all metrics, encompassing Progress Along Expert (Progress), Collisions, Speed Limit Compliance (Speed Limit), Drivable Area Compliance (Drivable), and TTC.
IV-C Ablation Study
TABLE II: Ablation Study
Remark IDD DS Horizon Score
No Additional Loss - - 81.47%
Decision Scope Absent 0 84.69%
Position 20 85.80%
Direct Velocity Loss - - 83.60%
Velocity 20 86.77%
Velocity 10 85.85%
Velocity 30 85.76%
Velocity 40 84.37%
To demonstrate the effectiveness of decision scope mechanism and IDD, ablation studies are conducted. The results are shown in Table II. In these studies, all models are trained on 20% of the dataset for 35 epochs and evaluated on the Random14 [25] test set, comprising 249 scenarios. The ablation study encompasses the following configurations: In the first experiment, we train the baseline model without other additional peripheral loss achieving a CLS-NR score of 81.47%. For comparison, we implement the IDD to the model to demonstrate the effectiveness of iterative decoder, which results in a CLS-NR score of 84.69%, a 3.22% improvement over the baseline. In the third experiment, we use the positional decision scope, enhancing the model’s CLS-NR score to 85.80%, a 4.33% increase from the baseline. In the fourth experiment, we introduce the direct velocity loss to the model, which yields a CLS-NR score of 83.60%, indicating a slight performance boost from emphasizing velocity differences. This result shows that emphasizing the velocity difference in demonstration can slightly improve the performance. In the experiment combining IDD with velocity decision scope, the model achieves a CLS-NR score of 86.77%, surpassing the baseline by 5.30% and the direct velocity loss by 3.17%. Further experiments vary the DS Horizon to explore the optimal effectiveness, and the results indicate that a DS Horizon of 20 achieves the best performance. Besides, with a total of 80 future time steps, a horizon of 40 represents level 0 detail data point count, effectively using all data for training. The DS mechanism exhibits peak performance at a horizon of 20, with a 2.40% decline in performance at a horizon of 40. This suggests that an overly extended planning scope can introduce excessive noise, and this confuses the model and degrades the planning performance, making the performance even worse than ignoring additional supervision.
@@ V Conclusions
This paper introduces a novel approach for planning future trajectories by extracting decisions across various frequency bands through wavelet decomposition, thereby enabling our model to discern the necessary decisions from log replays. We develop a mechanism by learning within the decision scope to accomplish this objective. References for multiple scope decisions are derived from the wavelet transformation of the expert trajectory data and utilized as an auxiliary loss for model training. The IDD technique recursively generates future trajectory details at multiple detail levels through the transformer. The experimental results indicate that our method outperforms baseline methods in achieving higher driving scores within the nuPlan dataset. Notably, in the small sample quantity comparison experiment, our method demonstrates a significant improvement of up to 5.77% in driving scores compared with the baseline. In the future, we plan to explore additional wavelet functions and potentially define a custom mother wavelet to more effectively capture driving information across different scales. Furthermore, we intend to integrate our method with perceptual information to optimize perception range, frequency, and resolution for key perceptual entities, thereby conserving computational resources and enhancing driving performance.

### v2

@@ III-D2 Iterative Detail Decoder (IDD)
$\frac{1}{N+1}$ $\left(\sum^{N}_{l=0}\left(\lVert\mathcal{T}_{D}^{l}[:H_{l}]-\hat{\mathcal{T}}_{D}^{l}[:H_{l}]\rVert_{2}\right)+\lVert\mathcal{T}_{A}^{N}-\hat{\mathcal{T}}_{A}^{N}\rVert_{2}\right).$
@@ IV Experiments
TABLE I: Preliminary Experiments
Detail Gen. Supervision
# Iterrative Detail heads Horizon DWT DWH Remarks CLS-NR
1 - - 80 - - PLUTO, m6 83.28%
2 - - 10 - - Truncation 80.12%
3 - - 20 - - Truncation 87.92%
4 - - 40 - - Truncation 85.22%
5 - - - - - Time decay, $l=e,p=1$ 88.36%
6 - - - - - Time norm 87.97%
7 - - - - - Time norm+contras 92.22%
8 - 10 - - 85.52%
9 - 10 - - 82.43%
10 - 10 - Position 82.26%
11 - 10 - Position 84.71%
12 - 20 - Position 82.48%
13 - 20 - Position 82.96%
@@ IV-A Implementation Details
Our training is carried out in the nuPlan dataset [2], which comprises a training set of 25,000 scenarios. We evaluate our model’s performance against baseline models within the nuPlan challenge framework. In addition, we conduct an ablation study to determine the significance of each component within our model. The training is performed on a server equipped with 8 NVIDIA L20 GPUs. We employ the Adam optimizer, with a start learning rate of 0.001, 3 warm-up epochs, and a batch size of 32 for each GPU node, over the course of 25 training epochs. The PyTorch library is utilized for the model’s implementation, with each training session approximately lasting 2 days.
@@
Our evaluation is conducted on the Val14 [23] subset of the nuPlan dataset, encompassing 1,090 validating scenarios. Each simulation entails a 15-second rollout at a frequency of 10 Hz. The evaluation score is derived using the official evaluation script provided by the nuPlan challenge. Our primary comparison metric is the CLS-NR driving score, which is more challenging as other vehicles do not avoid collisions with the ego vehicle. Besides, the future steps of our model is 80 while the history steps is 21. Both our model and PLUTO [19] are set to generate 6 modal trajectories. In our experiments, the maximum horizons of decision scope across levels are controlled by setting DS Horizon, which is denoted as $h$. Furthermore, we borrow the idea in [24] to facilitate the rapid implementation and execution of wavelet transformation in our experiments.
IV-B Preliminary Experiments
In order to identify truly effective strategies from the ones we explored, extensive comparative experiments are conducted. In these studies, all models are trained on 20% of the dataset for 35 epochs and evaluated on the Random14 [23] test set, comprising 249 scenarios. The results are shown in Table I. The ablation study encompasses the following configurations: In the first experiment, we train the baseline model without additional peripheral losses, achieving a CLS-NR score of 83.28%. Rows 1–4 are designed to explore the existence of the PlanScope problem. In these experiments, we apply a Time Truncation Horizon of 10, 20, 40, and 80 time steps, respectively. The corresponding results are 80.12%, 87.92%, 85.22%, and 83.28%. The initially rising and subsequently falling score suggests that a planning horizon of 20 is relatively optimal. This corroborates our hypothesis that short-term decisions can improve planning performance. However, overly myopic decisions may introduce other issues, such as collisions due to the absence of predictive planning. Results from rows 5–7 demonstrate that weighting the regression loss enhances planning performance in closed-loop simulations. Specifically, the Decreasing Certainty weight is denoted as timedecay, and Time-Dependent Normalization is denoted as timenorm. The demonstrated value of $p$ and $l$ timedecay achieves best performance in our preliminary grid search. Row 7 shows timenorm is of good compatibility with the contrastive loss. Comparisons in rows 8–11 indicate that the Multi-head Detail Decoder exhibits better compatibility with DWH than IDD, while IDD performs well in learning supervision signals obtained via DWT. Additionally, the comparison of rows 10–13 shows that the combination of IDD and DWT achieves relatively better performance when the DS Horizon is set to 20.
IV-C Comparison with Baselines
Figure 5: The qalitative comparison is conducted between the standard PLUTO and PlanScope (Scope) with time dependent normalization (TN), whose quantitative results listed in Table II fourth and third row from bottom respectively. In scenario S1, PLUTO fails to make a timely response to the impending collision in the short-term, but chooses to continue along the long-term decision, while Scope makes a timely yielding maneuver. In scenario S2, PLUTO adopts a conservative decision due to the high uncertainty of the future trajectory caused by interaction with the rear vehicle in a long time horizon, while Scope adopts human-like lane change decision.
TABLE II: Comparison with Baseline Methods
Planner Comfortness Progress w/o Collision in Speed Limit Drivable TTC CLS-NR CLS-R
Log-Replay 99.27% 98.99% 98.76% 96.47% 98.07% 94.40% 93.68% 81.24%
IDM 79.31% 86.16% 90.92% 97.33% 94.04% 83.49% 79.31% 79.31%
RasterModel 81.64% 80.60% 86.97% 98.03% 85.04% 81.46% 66.92% 64.66%
UrbanDriver 100.00% 80.83% 94.13% 91.58% 90.83% 80.28% 67.72% 64.87%
GameFormer 93.39% 89.04% 94.32% 98.67 94.87% 86.77% 82.95% 83.88%
PLUTO-m6 95.14% 87.13% 94.45% 97.87% 98.90% 90.92% 85.81% 76.45%
PlanScope-Ih10-DWT 95.96% 86.04% 95.37% 98.20% 98.44% 91.74% 85.99% 75.81%
PlanScope-Mh10-DWH 96.42% 88.72% 95.73% 98.17% 98.72% 92.66% 87.77% 75.39%
PlanScope-Mh20-DWT 97.89% 86.70% 93.72% 98.42% 98.81% 91.28% 85.80% 77.60%
PlanScope-Th20 99.27% 76.32% 95.50% 99.22% 99.36% 91.93% 83.86% 83.74%
PlanScope-timedecay 96.24% 86.34% 95.14% 98.24% 99.63% 91.56% 86.48% 75.22%
PlanScope-timenorm 98.26% 88.04% 96.28% 98.20% 99.45% 92.94% 88.41% 75.56%
PLUTO-m12-C 96.41% 93.28% 96.18% 98.13% 98.53% 93.28% 89.04% 80.01%
PlanScope-timenorm-m12-C 97.25% 89.54% 97.29% 98.04% 99.72% 93.85% 91.32% 80.96%
PLUTO-m12-C-H 91.93% 93.65% 98.30% 98.20% 99.72% 94.04% 93.21% 92.06%
PlanScope-timenorm-m12-C-H 93.85% 93.09% 98.35% 98.22% 99.82% 94.86% 93.59% 91.07%
The suffixes m6/m12 denote that the trajectory mode number is set to 6 or 12, respectively. Ih10 abbreviates IDD with a DS Horizon of 10, while Mh10 refers to MDD with a DS Horizon of 10. Th20 signifies the use of Time Truncation weights with a DS Horizon of 20. C denotes the application of contrastive learning. H represents a hybrid method combining learning and rule-based post-processing. CLS-R represent Reactive Closed Loop Simulation.
@@ IV-C Comparison with Baselines
We conduct a comparative analysis of our method against several standard approaches, the comparative results are presented in Table II. Log-Replay, which entails a straightforward replay of expert trajectories, serves as a baseline reference. This demonstrates the capability of our simulator to accurately assess the performance of planning algorithms. IDM [25] employs the Intelligent Driver Model to maintain safe distances from other agents and adhere to the reference lane. RasterModel, introduced in [2], is a CNN-based planning model. UrbanDriver [26] is a vectorized planning model leveraging PointNet based polyline encoders and transformer architecture. PLUTO [19], our primary baseline, achieves state-of-the-art performance in the nuPlan planning challenge in April 2024. To save computing resources, the multimodal trajectory mode number is set to m6 at first. By default, we have excluded the rule-based trajectory selector, thereby demonstrating that our new model enhances the performance of the neural planner. Besides, we also eliminate the contrastive learning module, which is empowered by rule-based positive and negative sample construction. These modules are eliminated to reduce artifacts that may introduce bias in results. Our model, designated as PlanScope, is evaluated with various modules performs well in preliminary experiments. Part of qualitative comparison is depicted and introduced in Fig. 5. The results in the table indicate that incorporating timenorm further enhances PLUTO’s performance in closed-loop simulations, achieving a total score of 93.59%. This improvement is primarily constrained by the capabilities of the post-processing module. Under a framework that relies solely on network-based output without post-processing, our closed-loop total score 91.32% increases by 2.28% compared to the previous result. In experiments with mode number equals 6 and without contrastive learning, the model with timenorm also achieves best CLS-NR score to 88.41%, which increases 2.6% compared to PLUTO-m6. We also observe that the score for Time Truncation is lower than the baseline, suggesting that a decision-making strategy focused solely on the short term may overlook critical long-term aspects. Additionally, methods based on detail generation and decomposition show slight improvement. More decomposition methods, decomposing measurements, wavelet forms and related research on horizon adaptation might be required to further enhance its performance. Furthermore, we also found that the CLS-R score improved with Time Truncation. The specific mechanism that lead to improvements on CLS-R score needs to be further explored.
@@ V Conclusions
This paper identifies the decision sequence learning problem and formulates it as PlanScope, which has been overlooked in previous closed-loop simulations. Several approaches for enabling model to extract necessary decisions from log replays are proposed and explored. Among these, the implementation of timenorm achieves a state-of-the-art 91.32% CLS-NR score in no-post-processing mode. In the hybrid mode, it reaches a 93.59% CLS-NR score, which also surpasses the baseline. In addition, detail generation modes combined with our proposed decomposition methods show limited improvement. Their performance does not surpass that of timenorm and timedecay. Nonetheless, wavelet transformation emerges as a potential solution to this issue. We present these findings as a starting point to narrow down the exploration range for future research. In future work, we plan to conduct more thorough experiments to explore the mechanisms underlying this problem, including further investigation of combinations and loss balancing across levels of detail.